OpenStudy (anonymous):

a person rolls two dice. what are the odds against throwing at least a nine?

jimthompson5910 (jim_thompson5910):

P(At least 9) = P(9) + P(10) + P(11) + P(12) P(At least 9) = 4/36 + 3/36 + 2/36 + 1/36 P(At least 9) = 10/36 P(At least 9) = 5/18 Since 18-5 = 13, this means that the odds against rolling at least a 9 are 13:5

OpenStudy (anonymous):

that isnt one of my answers though...could it be 12:5??

jimthompson5910 (jim_thompson5910):

what are your other answer choices?

OpenStudy (anonymous):

1:5, 5:1, 6:1, 1:6, 5:7, 7:5, 5:12, 12:5 or none of these

jimthompson5910 (jim_thompson5910):

None of other choices fit, so I would say "none of these". It's either that, or there's a typo.

OpenStudy (anonymous):

thank you!